OpenStudy (anonymous):

What is the solution set for |-4x - 4| - 28 = -4?

Parth (parthkohli):

1) add 28 to both sides. 2) the equation will look something like this - \(-4x - 28 = \pm24\)

OpenStudy (anonymous):

ok soo.. im going too add 28 too both sides of the equation . i got equation, -4x=52

Parth (parthkohli):

plus AND minus.

Parth (parthkohli):

\( \color{Black}{\Rightarrow -4x - 4 = + 24 }\) \( \color{Black}{\Rightarrow -4x - 4 = -24}\)

Parth (parthkohli):

I meant -4x - 4 in the first place..sorry

OpenStudy (earthcitizen):

\[4 (\left| x+1 \right| -7) = -4\] divide both side by 4 \[\left| x+1 \right| = 6\] Square both sides\[(x+1)^{2} = 36\] Expand and solve for x \[x^2+2 x+1 = 36\] \[x^2+2 x-35 = 0\] can you find x ?

Parth (parthkohli):

What?

Parth (parthkohli):

Oh I see. Another way to solve eh?

OpenStudy (earthcitizen):

what's the p @ParthKohli ?

Parth (parthkohli):

P?

OpenStudy (earthcitizen):

yh, problem. lol

Parth (parthkohli):

idk

Parth (parthkohli):

what do you mean by 'p'?

OpenStudy (earthcitizen):

yh, it's another way to solve the q

Parth (parthkohli):

:/

OpenStudy (earthcitizen):

ty